OpenStudy (anonymous):

Essay: Show all work. A gardener wants to create a rectangular garden with length of 9x - 5y ft. and width of x - 6y ft. What is an algebraic expression for the area of the garden? Be sure to multiply this out, express in simplest correct mathematical form, and include units.

OpenStudy (bahrom7893):

They make you write essays in math now?

OpenStudy (anonymous):

yeah they just want you to show all your work

OpenStudy (bahrom7893):

Essays.. in math...

OpenStudy (anonymous):

lol picture is to funny

OpenStudy (bahrom7893):

(9x-5y)*(x-6y) btw that's not me in the pic

OpenStudy (anonymous):

ok lol

OpenStudy (bahrom7893):

ohh include units.. anyway Area is length times width, so: (9x-5y)(ft)*(x-6y)(ft) = (9x-5y)(x-6y)(ft^2)

OpenStudy (anonymous):

is there more to it

OpenStudy (bahrom7893):

yea

OpenStudy (anonymous):

ok what comes after that

OpenStudy (bahrom7893):

multiply that whole thing out... (9x*x-9x*6y-5y*x+5y*6y)(ft^2) = = (9x^2 - 54xy + 30y^2)(ft^2)

OpenStudy (bahrom7893):

is that the simplest form anyone? i feel ancient, don't even know what simplest means.. maybe pull out a 3?

OpenStudy (anonymous):

lol is that the whole thing

OpenStudy (bahrom7893):

I think so.. i mean is the simplest form multiplied out or factored?

OpenStudy (anonymous):

ok thanks

OpenStudy (bahrom7893):

ACCESSDENIED!

OpenStudy (accessdenied):

You forgot your -5xy! (9x*x-9x*6y-[[[[[5y*x]]]]]+5y*6y)(ft^2) = (9x^2 - 54xy + 30y^2)(ft^2) Don't forget those pesky -5xy's!

OpenStudy (bahrom7893):

no i didn't.. did i?

OpenStudy (bahrom7893):

oh yea i did, but so did u lol.

OpenStudy (bahrom7893):

NATZ COME BACK!

OpenStudy (accessdenied):

Well, that was just quoting you. ;p It would be -59xy instead.

OpenStudy (bahrom7893):

lol kk

OpenStudy (anonymous):

i am back

OpenStudy (bahrom7893):

it was -59xy instead of -54xy

OpenStudy (anonymous):

ok so it is laid out like this (9x*x-9x*6y-[[[[[5y*x]]]]]+5y*6y)(ft^2) = (9x^2 - 54xy + 30y^2)(ft^2)

OpenStudy (accessdenied):

Sorry, that's just my pointing out the -5xy with lots of brackets. :p instead of "-54xy", you would have "-59xy." The rest is fine.

OpenStudy (anonymous):

oh ok got it thank u so much